OpenStudy (leann01xxx):

It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? (Round answer to nearest tenth.)

OpenStudy (hayleymeyer):

Hmmmm

OpenStudy (hayleymeyer):

Let d = dist traveled one way Write a time equation; time = dist/speed d%2F20 + d%2F30 = 12 multiply eq by 60, results 3d + 2d = 60(12) 5d = 720 d = 720%2F5 = 144 mi one way : Find the time at each speed 144%2F20 = 7.2 hrs at 20 mph 144%2F30 = 4.8 hrs at 30 mph

OpenStudy (hayleymeyer):

Does that make since at all? >_<

OpenStudy (leann01xxx):

A little but not really :/